OpenStudy (anonymous):

Which line passes through the point (10,-1) and has a slope of -2?

OpenStudy (anonymous):

a line passing through a given point (x,y) with a given slope m is given by;- (y-y1)=m(x-x1)

OpenStudy (anonymous):

here the given point is (10,-1) and slope is -2 so the required equation is (y-(-1))=-2(x-10) y+1=-2x+20 y+2x=19

OpenStudy (anonymous):

to get y alone would it just be y=-2x+19?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome!